OpenStudy (anonymous):

Find the x-intercepts of the graph y=x^2+8x+12?

OpenStudy (anonymous):

let y=0 (x+2(x+6) =0 x= -2 x=-6

OpenStudy (anonymous):

to find the x intercepts set y = 0 0 = x² + 8x + 12 now solve for x, best way to do it is factoring 0 = (x + 6)(x + 2) now separate them and solve for x x + 6 = 0 x = -6 x + 2 = 0 x = -2 the x intercepts are (-2,0) and (-6,0)